OpenStudy (anonymous):

write an equation of the line. (0,3) 7x+3y=2

OpenStudy (mathstudent55):

I don't understand the question. (0, 3) is a point. 7x + 3y = 2 is the equation of a line. What exactly is the problem asking for?

OpenStudy (anonymous):

it says....write an equation of the line containing the given point and perpendicular to the given line.

OpenStudy (anonymous):

@mathstudent55

hartnn (hartnn):

can you find the slope of the line 7x+3y = 2 ?

OpenStudy (anonymous):

-7/3

hartnn (hartnn):

thats correct! did you know that product of slope of perpendicular lines = -1 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

now i do lol

hartnn (hartnn):

now you know m*(-7/3) = -1 find m from here :)

OpenStudy (anonymous):

wait huh? i dont understand.

hartnn (hartnn):

say the slope is a/b then the slope of perpendicular line will be -b/a or "the negative reciprocal"

OpenStudy (anonymous):

so it would be 7/3=-1

hartnn (hartnn):

the slope of line was -7/3 so the slope of perpendicular line will be +3/7 i interchanged the numerator and denominator (reciprocal) and inverted the sign

OpenStudy (anonymous):

so how would i find the equation of the line?

OpenStudy (anonymous):

y=

hartnn (hartnn):

now you have the slope m = 3/7 and a point x1,y1 = 0,3 just plug these in \(y-y_1 = m (x-x_1)\)

OpenStudy (anonymous):

y-3=3/7(x-0)

hartnn (hartnn):

correct! now simplify

OpenStudy (anonymous):

y=3x/7+3

hartnn (hartnn):

y-3= 3x/7 multiply 7 on both sides first

OpenStudy (anonymous):

21-3=3x/7

hartnn (hartnn):

7(y-3) =3x 7y -21 = 3x 7y-3x = 21

OpenStudy (anonymous):

so how would i turn that into y=mx+b format?

hartnn (hartnn):

oh, u need it into y=mx+b format ? then y =(3/7)x+3 was correct :)

OpenStudy (anonymous):

do i usethe parenthesis?

hartnn (hartnn):

yes, you can but both y= (3/7)x+3 and y= 3x/7 + 3 are correct. though the second on is bit confusing